OpenStudy (anonymous):

What is x? x * sqrt(3)=24

OpenStudy (bmk614):

x=24/sqr(3)

OpenStudy (phi):

divide both sides by sqr(3) most people would "rationalize" the answer, by multiplying top and bottom by sqr(3)

OpenStudy (anonymous):

@phi I get 24*sqrt(3). When you multiply that by the sqrt(3) you get 72..

OpenStudy (anonymous):

This is for a 30-60-90 triangle #8

OpenStudy (phi):

when you have for example 2x and you divide by 2 \[ \frac{2x}{2} \] that is the same as \[ \frac{2}{2} x\] and that is 1*x or just x in other words, anything divided by itself is 1

OpenStudy (phi):

so if you have \[ x \sqrt{3} = 24\] you divide both sides by sqr(3) \[ \frac{\sqrt{3}}{\sqrt{3}} x = \frac{24}{\sqrt{3}} \]

OpenStudy (anonymous):

And

OpenStudy (anonymous):

@phi but then you rationalize it and get 24*sqrt(3)

OpenStudy (phi):

so you are doing Q8 for a 30-60-90 ? they give the "long side" across from the 60 deg it is sqr(3) bigger than the "short side" so that is why you did x sqr(3)= 24 where x is the length of side YZ

OpenStudy (phi):

after dividing by sqr(3) we have \[ x = \frac{24}{\sqrt{3}} \] now rationalize by multiplying top and bottom by sqr(3) \[ x = \frac{24}{\sqrt{3}} \cdot \frac{\sqrt{3}}{\sqrt{3}} \]

OpenStudy (phi):

the top is sqr(3) * 24 the bottom is 3

OpenStudy (phi):

you get \[ x = \frac{24\sqrt{3}}{3} = 8 \sqrt{3} \]

OpenStudy (anonymous):

@phi x is the length of yz and 2x is xy and xz is x*sqrt(3)

OpenStudy (phi):

that is the short side (opposite the 30 deg angle) the hypotenuse will be twice that, or 16 sqr(3)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@phi what is 15sqrt(3)/3?

OpenStudy (phi):

you can thing of it as 15 * sqr(3) * ⅓ and rearrange the order (right?) to ⅓ * 15 * sqr(3) what is ⅓ * 15 ?

OpenStudy (anonymous):

@phi this is or #10 for a 30-60-90. 5?? That doesn't seem right

OpenStudy (phi):

you should get \[ 5 \sqrt{3} \]

OpenStudy (phi):

and the hypotenuse will be 10 sqr(3)

OpenStudy (anonymous):

Oh ok @phi thank you again